OpenStudy (eric_d):

Expressing x^3+1/x^2+x-2 in partial fraction

OpenStudy (eric_d):

Quotient=x-1 Remainder=3x-1

OpenStudy (eric_d):

x^3+1/(x-1)(x+2)=(x-1)+3x-1/(x-1)(x+2) =(x-1)+A/x-1+B/x+2 X^3+1=(x-1)^2(x+2)+A(x+2)+B(x-1) When x=1, A=2/3 When x=-2, B=7/3 Thus, x^3+1/x^2+x-2= x-1+ 2/3(x-1)+7/3(x+2)

OpenStudy (eric_d):



OpenStudy (eric_d):

Can those two be the final answer ?

OpenStudy (eric_d):

@amistre64

OpenStudy (amistre64):

x^3+1 ------- x^2+x-2 can we factor them? (x+1)(x^2-x+1) --------------- (x+2) (x-1) hmm, not unless you had a typo long division then ..... assuming youve covered this x - 1 ---------------- x^2+x-2 | x^3+1 -(x^3+x^2-2x) --------------- -x^2+2x+1 -(-x^2 -x+2) -------------- 3x-1 is our remainder

OpenStudy (amistre64):

partial fraction decomposition .... is prolly what your refering to :/

OpenStudy (eric_d):

yes

OpenStudy (amistre64):

so factored form is great lol (x+1)(x^2-x+1) A B --------------- = ------- + ----- (x+2) (x-1) x+2 x-1 when x=-2 we can solve for A with coverup (-2+1)(4+2+1) --------------- = A = 7/3 (-2-1) when x=1 we can solve for B with coverup (1+1)(1-1+1) --------------- = B = 2/3 (1+2) assuming ive done it correctly lol

OpenStudy (eric_d):

okay

OpenStudy (amistre64):

if we simply work the remainder 3x-1 --------- (x+2)(x-1) 3(-2)-1 --------- = A = 7/3 (-2-1) 3(1)-1 --------- = B = 2/3 all the same (1+2)

OpenStudy (eric_d):

Okay... I'm confused with this.... Express Partial fraction 2x^2+3x+2/x^2+3x+2

OpenStudy (eric_d):

The quotient is 2 Remainder is -3x-2

OpenStudy (amistre64):

and the factors of the denominator?

OpenStudy (eric_d):

Nxt, I need to simplify rite

OpenStudy (eric_d):

x^2/(x+1)(x+2)= A/(x+1)+ B(x+2) x^2/(x+2)(x+1)=A(x+2) + B(x+1) Are these considered the same Will it affect the final answer?

OpenStudy (amistre64):

those are not considered the same

OpenStudy (amistre64):

\[\frac{-3x-2}{(x+2)(x+1)}=\frac A{x+2}+\frac B{x+1}\] now lets mutliply thru by (x+2)(x+1) \[\bcancel{\cancel{\frac{(x+2)(x+1)}{(x+2)(x+1)}}}(-3x-2)=\frac {A\cancel{(x+2)}(x+1)}{\cancel{x+2}}+\frac {B(x+2)\cancel{(x+1)}}{\cancel{x+1}}\] \[-3x-2=A(x+1)+B(x+2)\]

OpenStudy (eric_d):

Based on this working... -3x-2/(x+2)(x+1)= A/x+2 + B/x+1 What if I write like this..... -3x-2/(x+1)(x+2)=A/x+1 + B/x+2 Is this considered the same?

OpenStudy (amistre64):

its the same generality, but you wont get the same A=A and B=B results

OpenStudy (amistre64):

A and B are just some unknowns that you are working to find. their names are immaterial really. But if your expecting that the name A has some magical mystical property that it must maintain in all expressions then your thought process would be a little off

OpenStudy (eric_d):

So, if I swith them, the final answer might be different?

OpenStudy (amistre64):

it wont be 'different'. As long as your working it consistently

OpenStudy (eric_d):

Ok, thanks Amistre!

OpenStudy (amistre64):

youre welcome.